OpenStudy (anonymous):

find the equation of the normal line to the curve y=ln(x^2-3) at x=2 (step by step) WILL FAN AND MEDAL!

OpenStudy (amistre64):

first find the tangent, and also the value of f(2)

OpenStudy (amistre64):

take the derivative .... y' defines the slope of the tangent at a given point

OpenStudy (anonymous):

would the derivative be ln(2x)??? @amistre64

OpenStudy (amistre64):

the derivative of ln(u) = u'/u let u = x^2 -3, so u' = 2x

OpenStudy (amistre64):

y' = 2x/(x^2-3) .... and at x=2 this is?

OpenStudy (anonymous):

i plug the 2 into the x's right? if so it would be -2 @amistre64

OpenStudy (amistre64):

x=2 2(2)/(2^2-3) = 4/(4-3) = 4

OpenStudy (amistre64):

not real sure how youd get -2 .... but thats beside the point :)

OpenStudy (anonymous):

yea......i dont know either but thanks!! @amistre64

OpenStudy (amistre64):

the important thing is knowing the slope of the tangent line allows us to get the slope of the normal line by perping it ... tangent and normal are 90 degrees to each other meaning if the tangent is m, the normal is -1/m

OpenStudy (amistre64):

now tell me what you know about contructing a line using a point and a slope

OpenStudy (anonymous):

use rise/run (slope) when the points are plotted on the graph to make a line

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

sites lagging ... lets go this route since you gave it a shot given slope m, and a point (a,b) the equation of a line in point slope form is constructed as: y-b = m(x-a)

OpenStudy (amistre64):

in terms of functions and their derivatives .. our tangent line is: \[y-f(a)=f'(a)(x-a)\] our normal line is: \[y-f(a)=-\frac{1}{f'(a)}(x-a)\]

OpenStudy (anonymous):

what is f(a) and f'(a) im confused now??? @amistre64